Evaluate ((-2)^2-2)^2+5 *4
Found 3 solutions by palanisamy, mananth, ikleyn:
Answer by palanisamy(496)

((-2)^2-2)^2+5 *4 = [4-2]^2+20
= 2^2+20
= 4+20
= 24

Answer by mananth(16949)

Evaluate ((-2)^2-2)^2+5 *4

((+4)-2))^2+5
(+2)^2+5 *4
4+5 *9
9*9=81

Answer by ikleyn(53996)

.
Evaluate ((-2)^2-2)^2+5*4
~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        What is written in the post by @mananth is absurdist.
        See my correct solution below.



((-2)^2-2)^2 + 5*4 = (4-2)^2 + 5*4 = 2^2 + 5*4 = 4 + 20 = 24.         <<<---===     ANSWER

Solved correctly.